THE OBSERVATION AND ANALYSIS OF THE PURE ROTATIONAL SPECTRA OF X0^{+} PtS AND X0^{+} WO

\underline{S. A. COOKE}, C. KRUMREY, D. K. RUSSELL^{a} and M. C. L. GERRY, Department of Chemistry, The University of British Columbia, 2036 Main Mall, Vancouver, British Columbia, Canada, V6T 1Z1.

Measurements of the pure rotational spectra of 8 isotopomers of PtS using Fourier transform microwave (FTMW) spectroscopy have been reported previously^{a}. Pure rotational transitions from 8 isotopomers of WO have been recorded for the first time using FTMW spectroscopy. WO shares the same electronic ground state as PtS, X0^{+}, and modified Dunham-type spectral analyses of the data sets have been performed in parallel for the two molecules. For the analyses abnormally large Born-Oppenheimer breakdown (BOB) terms were required for all of the atoms in order to fit the data to within the experimental uncertainties. The magnitude of these BOB terms will be discussed along with the influence of field-shift effects arising from the Pt and W nuclei on the rotational energy levels. Nuclear spin-rotation constants for 195Pt in PtS and ^{183}W in WO have also been recorded.

aOn sabbatical from Department of Chemistry, The University of Auckland, Private Bag 92019, Auckland, New Zealand

aS. A. Cooke and M. C. L. Gerry, Paper TJ10, 57^{th} International Symposium on Molecular Spectroscopy, Ohio State University,